OpenStudy (ja1):

I a trying to find a GCF to factor out this quadratic equation, can you help? 0 = -x^2 + 2x + 15

OpenStudy (ja1):

@mathstudent55

OpenStudy (espex):

What is something that will go into every term in your equation?

OpenStudy (ja1):

Excuse my mentality, I have been doing these right for the past half hour but I guess my mind decided to troll me on this question *facepalm*

OpenStudy (ja1):

Thanks for bringing me to senses, also would it be -x because I have never seen -x before :/ @eSpeX

OpenStudy (espex):

You are on the right track and I apologize if I came across as condescending. Your last term does not have a 'x' in it though...

OpenStudy (ja1):

No your fine :), oh and the way the lesson taught me is to just do this when there is no GCF: (x +/- #)(x +/- #)

OpenStudy (ja1):

Wait, grrr I am wrong, I am trying to find the GCF of 2 and 15 my bad

OpenStudy (espex):

Rather than try a -x, which as we said does not exist in every term, what about -1?

OpenStudy (espex):

GCF or are you factoring?

OpenStudy (ja1):

Yes but then the product would be negative and so it would be wrong right? Since 15 is positive and so is 2

OpenStudy (ja1):

I need to find two numbers that the sum is 2 and the product is 15

OpenStudy (espex):

No, you are correct, and so what multiplies together to create -15 and adds together to get -2?

OpenStudy (ja1):

The original equation is: x^2 = 2x + 15

OpenStudy (ja1):

Well in that case it would be +5 -3

OpenStudy (ja1):

nono -5 +3

OpenStudy (espex):

I would have pushed everything to the left as to keep my square positive. And +5-3=+2

OpenStudy (espex):

Exactly. Good job. So your final result is?

OpenStudy (ja1):

(x + 3)(x - 5) x + 3 = 0 OR x - 5 = 0 -3 -3 +5 +5 x = -3 AND x = 5 right?

OpenStudy (espex):

Yes, that is right. I would point out though that what you have done is considered "factoring" and not finding the GCF. :)

OpenStudy (ja1):

Yes me and correct terminology do not get along xD but I usually understand it that way in my head but thanks :D

OpenStudy (espex):

You're welcome. Best of luck.

OpenStudy (ja1):

Thanks.